OpenStudy (anonymous):

Write an equation of a line in slope intercept form with slope 1/3 going through the point (-6, 2). Show your work

OpenStudy (anonymous):

slope intercept form : y = mx + b using slope 1/3 and points (-6,2) 2 = 1/3(-6) + b can you solve for b ?

OpenStudy (anonymous):

no idea... v~v

OpenStudy (caozeyuan):

y-2=1/3)*(x+6 so y=1/3)*x+4

OpenStudy (anonymous):

2 = 1/3(-6) + b 2 = -2 + b 2 + 2 = b 4 = b Therefore, you equation is : y = 1/3x + 4

OpenStudy (anonymous):

thankyou sooooooo much \(^~^)/

OpenStudy (anonymous):

your very welcome :)

OpenStudy (anonymous):

wat about with points (2, 1) and (- 1, - 5). o.o

OpenStudy (anonymous):

is this a new problem or part of the last problem ?

OpenStudy (anonymous):

not at all http://openstudy.com/updates/51dc0d9fe4b076f7da403af6